John is one year less than three times as old as Ann. the sum of their ages is 79. What are their ages?
Answer by ikleyn(52788)

.
From the condition, you have these two equations


    J = 3A - 1     (1)     ("John is one year less than three times as old as Ann.")

    J + A = 79.    (2)     (It is clear without explanations)


From eq(1), substitute the expression for J into eq(2). You will get


    (3A-1) + A = 79

    4A = 79 + 1

    4A = 80  ====>  A = 80/4 = 20.


Answer.  Ann is 20 years old.  John is 79-20 = 59 years old.

Solved.